OpenStudy (anonymous):

(-e^-k) - (-e^6) = 0.5 Need to find k. I got to: -e^-k = 0.4976 How do I go from here to free k and find it out? Thank you in advance.

OpenStudy (p0sitr0n):

take the logarithm on both sides -k*ln(e)=ln(e^6-.5) k=-ln(e^6-5)